OpenStudy (meilendurcer):

Shayla is baking muffins that require 1/2 cup vegetable oil. She only has 1/3 cup of vegetable oil left. How much more does she need? write out

OpenStudy (anonymous):

Find a common denominator.

OpenStudy (valpey):

\[\frac{1}{2}-\frac{1}{3}\]The least common multiple of 2 and 3 is 6, so we multiply each fraction so that the denominator is 6:\[\frac{1*3}{2*3}-\frac{1*2}{3*2} =\frac{3}{6}-\frac{2}{6}\]

OpenStudy (meilendurcer):

wow. thanks!!